# PhysicsCorrect: A Training-Free Approach for Stable Neural PDE Simulations

### v1

@@ Results.
Figure 5 shows the long-term rollout performance of different neural architectures with and without our correction framework. Even with the improved second-order formulation, baseline models (particularly ViT and UNet) still exhibit error growth over time. Our physics-informed corrector consistently enhances prediction accuracy across all architectures, maintaining low PDE residuals throughout the simulation.
The wave equation results demonstrate our method’s effectiveness on linear PDEs with oscillatory dynamics. Interestingly, the performance improvement is more pronounced for architectures that struggle more with the baseline formulation (ViT and UNet), suggesting that our correction approach can help compensate for architecture-specific weaknesses in capturing certain physical dynamics.
@@ 4.3 Kuramoto-Sivashinsky equation
Our final and most challenging test case is the Kuramoto–Sivashinsky (KS) equation – a fourth-order nonlinear PDE that exhibits chaotic dynamics (Appendix B.3). This system tests our method’s effectiveness on strongly nonlinear, chaotic dynamics where prediction errors can amplify rapidly and where linearization approximations are most severely challenged.
@@
We simulate the KS equation on a spatial domain $[0,64]$ with a resolution of 512 points, focusing on the chaotic regime. Starting from $v(x,50)$, we generate 512 trajectories for training (first 500 steps) and 64 trajectories for testing (full 1000 steps) using a spectral method with temporal step size 0.05 (dresdner_learning_2023, ).
@@
Surprisingly, as shown in Figure 5, our approach with cached pseudoinverse still provides significant stability improvements despite using an approximation of the true Jacobian. Figure 7(a) shows a qualitative comparison between baseline FNO predictions (top) and corrected predictions (bottom), demonstrating that our method successfully maintains the complex spatiotemporal patterns of the KS equation over long rollouts.
To further investigate the impact of Jacobian approximation in chaotic systems, we conducted additional experiments varying the frequency of Jacobian updates, as shown in Figure 7(b). The results reveal that recalculating the Jacobian every 3-10 time steps provides minimal improvement over the fully cached approach (updating only once at initialization). This suggests that for the KS equation, the primary benefit comes from the initial projection toward the physically consistent manifold rather than from having a perfectly accurate Jacobian at each step.
These findings lead to an important insight: accurate definition of the PDE residual is more critical than perfect estimation of the Jacobian pseudoinverse. Even with an approximate linear correction, projecting predictions toward the physically consistent manifold provides sufficient regularization to prevent error accumulation. This observation is particularly significant for chaotic systems, where traditional methods often struggle to maintain long-term stability, and offers a favorable trade-off between computational efficiency and prediction accuracy.
@@ 5 Discussion
Despite its effectiveness, PhysicsCorrect faces three key limitations that suggest directions for future work: (1) computational scalability to high-resolution simulations, as Figure 8 demonstrates quadratic scaling of time and memory requirements with resolution, which could be mitigated by inverting the Jacobian approximately using iterative methods that only require Jacobian-vector products; (2) discretization-induced errors creating an inherent gap between our correction target and true numerical references, potentially improvable through higher-order discretization schemes; and (3) potential breakdown of linearization approximations in extreme chaotic systems, which might be addressed by hybrid approaches combining efficient linear correction with occasional nonlinear optimization steps for systems with extreme sensitivity or strong nonlinearities.
Figure 8: PhysicsCorrect’s computational scaling with grid resolution for 2D Navier-Stokes. Plots show computation time (left), GPU memory usage (center), and peak memory consumption (right). All metrics scale quadratically with resolution, highlighting memory and compute requirements as the primary limitation for high-resolution applications.
@@ Appendix A PDE Residuals Approximation
We detail the numerical discretization employed in this work for defining the PDE residuals, along with relevant implementation considerations. In all experiments, the computational domain $\Omega\subseteq\mathbb{R}^{2}$ is discretized into an $n\times n$ uniform Cartesian grid with mesh spacing $\Delta$. Any function defined on $\Omega$ is then approximated by its values at the grid points. Specifically, we denote the discrete approximation of a function function $u(x,y)$ at the gird point $(x_{i},y_{j})$ by $u_{i,j}$, where $x_{i}=x_{0}+i\Delta$ and $y_{j}=y_{0}+j\Delta$ for $i=0,\cdots,n-1$ and $j=0,\cdots,n-1$. For time-dependent variables, we discretize the time interval $[0,T]$ with a uniform time step $\Delta t$, and denote a discrete function $u$ at the spatial gird point $(i,j)$ and $t$-th time step by $u_{i,j}^{t}$.
2D Navier-Stokes equation. Considering incompressible two-dimensional Navier-Stokes equations with periodic boundary conditions as follows:
@@
$+0.5\frac{1}{\text{Re}\Delta^{2}}\left(\omega_{i+1,j}^{t}+\omega_{i-1,j}^{t}+\omega_{i,j+1}^{t}+\omega_{i,j-1}^{t}-4\omega_{i,j}^{t}\right.$
@@
2D Wave Equation. The two-dimensional wave equation with constant wave speed is formulated as
@@
at each grid point $(x_{i},y_{j})$ for $i=1,\cdots,n$ and $j=1,\cdots,n$. Since we did not have fine sampling in the temporal direction, we use an implicit scheme and central finite difference for the second-order time derivative. Given the displacement function $u^{t-1}_{i,j}$ at time $t-1$ and $u^{t}_{i,j}$ at time $t$, and the one-step prediction $\hat{u}^{t+1}_{i,j}$ at time $t+1$, the discretized PDE residual is represented as:
@@
Hence, $\frac{\partial L_{\text{PDE}}}{\partial\hat{u}^{t+1}}$ does not depend on the prediciton $\hat{u}^{t+1}$ itself. Notably, for the wave equation, being a linear system, the PDE residual in Equation 5 does not rely on the assumption of small prediction error, as no Taylor expansion is needed. This further implies that PhysicsCorrect remains robust to prediction errors in linear systems, a phenomenon that will be demonstrated empirically in later sections.
Kuramoto-Sivashinsky Equation. The Kuramoto-Sivashinsky equation with a constant viscosity parameter of 1.0 is as follows:
@@ Appendix B Experimental Details
In this section, we introduce the details of the configurations of different benchmarks and neural network configurations, as well as additional visualization results. The code and data will be available at https://github.com/summerwine668/PhysicsCorrect.
@@ B.1 Navier-Stokes Equation
Regarding the neural network architectures, we implement the standard Fourier Neural Operator (FNO) following [29], which consists of four Fourier layers. Each layer uses 12 Fourier modes in both the $x$ and $z$ directions, and the width of the feature map is set to 64. The activation function used is the Gaussian Error Linear Unit (GELU) [30]. After the Fourier layers, a linear projection is applied to map the high-dimensional latent representation back to the spatial domain. For the UNet architecture, we adopt the configuration described in [31], with a hidden size of 128. For the Vision Transformer (ViT), we use a patch size of 8 for spatial embedding and stack 4 standard pre-norm transformer blocks. Each block contains four attention heads, with each head having a dimensionality of 128. GELU is also used as the activation function for ViT.
@@
We fixed the wave speed $c$ as 1.0 in our test. We generate Gaussian random fields as initial conditions and then use fourth-order central differences for spatial derivatives and a fourth-order Runge–Kutta (RK4) scheme for time integration, to accurately evolve the displacement field, following [32]. The data are generated on a 128$\times$128 grid with a time step of 1e-4 and recorded every 100 time steps. We sampled 512 trajectories with the first 10 steps (with an interval of 1e-2) for training and another 64 trajectories with all 100 steps for testing. We use an initial learning rate of 1e-3 with a decay factor of 0.6 every 100 epochs.
@@ B.3 Kuranmoto-Sivashinksy Equation
We aim to solve the Equation 12 on a temporal domain $[0,100]$ and a spatial domain $[0,64]$ while focusing on the chaotic part. The equation is numerically solved using the spectral method [28], with a spatial interval of 0.125 and a temporal interval of 0.05, for data generation. We begin from $v(x,50)$ and train a one-step predictor using neural networks to learn the residual $\delta v^{t}=v^{t+1}-v^{t}$. The training set includes 512 trajectories, using only the first 500 time steps of each, while the test set contains 64 additional trajectories, each with the full 1,000-step rollout. All models are trained with a learning rate of 1e-3, a batch size of 256, and a total of 3,000 training epochs.
For the network configurations, we use a one-dimensional variant of the FNO (FNO-1D), adapted from its 2D counterpart by replacing all 2D convolutional operations with 1D convolutions. The number of Fourier modes is set to 16, and the feature map width is 64. GELU is used as the activation function. The 1D UNet is similarly adapted from the 2D UNet by substituting all 2D convolutional layers with 1D versions. The hidden size is set to 64. We use the LeakyReLU activation function with a negative slope of 0.1, along with Group Normalization [33]. For the 1D ViT, we use a patch size of 4 and an embedding dimension of 768. The model consists of 4 transformer blocks, each containing 16 attention heads with a head dimension of 768. GELU is employed as the activation function throughout the model.
@@
B.4 Additional Implementation details
Compute Resources. All training is performed on a single NVIDIA A6000 GPU. For the Navier-Stokes equation, training for 30,000 steps with a batch size of 64 takes roughly 60, 12, and 170 minutes using FNO, U-Net, and ViT, respectively. For the wave equation, training for 1,000 epochs takes roughly 29, 6, and 65 minutes using FNO, U-Net, and ViT, respectively. For the KS equation, training for 3,000 epochs takes roughly 100 minutes using FNO, 70 minutes using U-Net, and approximately 5 hours using ViT.
Negligible Additional Inference Time. We also evaluate the inference time of both the baseline models and their counterparts augmented with our proposed correctors. As shown in Figure 5, the ratio between the inference time of the baseline with corrector and that of the baseline model is close to 1.0. Although the Navier–Stokes case with the ViT baseline shows a slightly higher ratio, due to the lightweight nature of the baseline, the additional inference time introduced by our corrector remains negligible, particularly for larger-scale problems and more complex neural networks.
B.5 Sensitivity Analysis: Error Structure and Magnitude Limits

### v2

@@ Results.
Figure 4 shows the long-term rollout performance of different neural architectures with and without our correction framework. Even with the improved second-order formulation, baseline models (particularly ViT and UNet) still exhibit error growth over time. Our physics-informed corrector consistently enhances prediction accuracy across all architectures, maintaining low PDE residuals throughout the simulation. The wave equation results demonstrate our method’s effectiveness on linear PDEs with oscillatory dynamics. Interestingly, the performance improvement is more pronounced for architectures that struggle more with the baseline formulation (ViT and UNet), suggesting that our correction approach can help compensate for architecture-specific weaknesses in capturing certain physical dynamics.
@@ 4.3 Kuramoto-Sivashinsky equation
Our final and most challenging test case is the Kuramoto–Sivashinsky (KS) equation, a fourth-order nonlinear PDE that exhibits chaotic dynamics (more details are in the Appendix). This system tests our method’s effectiveness on strongly nonlinear, chaotic dynamics where prediction errors can amplify rapidly and where linearization approximations are most severely challenged.
@@
We simulate the KS equation on a spatial domain $[0,64]$ with a resolution of 512 points, focusing on the chaotic regime. Starting from $v(x,50)$, we generate 512 trajectories for training (first 500 steps) and 64 trajectories for testing (full 1000 steps) using a spectral method with temporal step size 0.05 (Dresdner et al. 2023).
@@
Surprisingly, as shown in Figure 4, our approach with cached pseudoinverse still provides significant stability improvements despite using an approximation of the true Jacobian. Figure 7(a) shows a qualitative comparison between baseline FNO predictions (top) and corrected predictions (bottom), demonstrating that our method successfully maintains the complex spatiotemporal patterns of the KS equation over long rollouts. To further investigate the impact of Jacobian approximation in chaotic systems, we conducted additional experiments varying the frequency of Jacobian updates, as shown in Figure 7(b). The results reveal that recalculating the Jacobian every 3-10 time steps provides minimal improvement over the fully cached approach (updating only once at initialization). This suggests that for the KS equation, the primary benefit comes from the initial projection toward the physically consistent manifold rather than from having a perfectly accurate Jacobian at each step.
These findings lead to an important insight: accurate definition of the PDE residual is more critical than perfect estimation of the Jacobian pseudoinverse. This, in turn, supports the validity of separating the residual evaluation from the Jacobian precomputation, as the correction remains effective even when the precomputed Jacobian is only approximate. Even with an approximate linear correction, projecting predictions toward the physically consistent manifold provides sufficient regularization to prevent error accumulation. This observation is particularly significant for chaotic systems, where traditional methods often struggle to maintain long-term stability, and offers a favorable trade-off between computational efficiency and prediction accuracy.
@@ 5 Discussion
Figure 8: PhysicsCorrect’s computational scaling with grid resolution for 2D Navier-Stokes. Plots show computation time (left), GPU memory usage (center), and peak memory consumption (right). All metrics scale quadratically with resolution, highlighting memory and compute requirements as the primary limitation for high-resolution applications, and indicating challenges in scaling to fully 3D problems.
@@
Despite its effectiveness, PhysicsCorrect faces three key limitations that suggest directions for future work: (1) computational scalability to high-resolution simulations, as Figure 8 demonstrates quadratic scaling of time and memory requirements with resolution, which could be mitigated by inverting the Jacobian approximately using iterative methods that only require Jacobian-vector products; (2) discretization-induced errors creating an inherent gap between our correction target and true numerical references, potentially improvable through higher-order discretization schemes; moreover, it can be naturally combined with spatio-temporal decomposition training frameworks (Huang et al. 2023), preserving the large spatial- and temporal-interval advantages of purely neural PDE solvers while enhancing stability through physics-based correction; and (3) potential breakdown of linearization approximations in extreme chaotic systems, which might be addressed by hybrid approaches combining efficient linear correction with occasional nonlinear optimization steps.
@@ Conclusions.
Acknowledgments
This work was supported by the U.S. Department of Energy, Advanced Scientific Computing Research program, under the “Resolution-invariant deep learning for accelerated propagation of epistemic and aleatory uncertainty in multi-scale energy storage systems, and beyond” project (Project No. 81824), and Penn AIxScience fellowship.
@@ Appendix A PDE Residuals Approximation
Regarding a general PDE in the form (Equation 1), it can be further decomposed as
$\mathcal{M}(\mathbf{u}_{t},\mathbf{u}_{tt})=\mathcal{L}(\mathbf{u})+\mathcal{N}(\mathbf{u}),$ (7)
where $\mathcal{L}$ denotes the linear term (e.g., diffusion, wave operator) and $\mathcal{N}$ denotes the non-linear term (e.g., advection). To approximate $\frac{\partial L_{\text{PDE}}}{\partial\hat{\mathbf{u}}_{t+1}}$ and enable the proposed caching strategy for efficiency, we generally treat linear terms in the PDE implicitly to ensure stability, while handling nonlinear terms explicitly to keep the Jacobian constant across timesteps. Under this semi-implicit scheme, the approximate Jacobian matrix becomes $\frac{I}{\Delta t}$-$\mathcal{L}$ for first-order equations or $\frac{I}{\Delta t^{2}}$-$\mathcal{L}$ for second-order equations.
We detail the numerical discretization employed in this work for defining the PDE residuals, along with relevant implementation considerations. In all experiments, the computational domain $\Omega\subseteq\mathbb{R}^{2}$ is discretized into an $n\times n$ uniform Cartesian grid with mesh spacing $\Delta$. Any function defined on $\Omega$ is then approximated by its values at the grid points. Specifically, we denote the discrete approximation of a function $u(x,y)$ at the grid point $(x_{i},y_{j})$ by $u_{i,j}$, where $x_{i}=x_{0}+i\Delta$ and $y_{j}=y_{0}+j\Delta$ for $i=0,\cdots,n-1$ and $j=0,\cdots,n-1$. For time-dependent variables, we discretize the time interval $[0,T]$ with a uniform time step $\Delta t$, and denote a discrete function $u$ at the spatial grid point $(i,j)$ and $t$-th time step by $u_{i,j}^{t}$.
2D Navier-Stokes equation.
Considering incompressible two-dimensional Navier-Stokes equations with periodic boundary conditions as follows:
@@
[…]
@@
2D Wave Equation.
The two-dimensional wave equation with constant wave speed is formulated as
@@
at each grid point $(x_{i},y_{j})$ for $i=1,\cdots,n$ and $j=1,\cdots,n$. Since we did not have fine sampling in the temporal direction, we use an implicit scheme and central finite difference for the second-order time derivative. Given the displacement function $u^{t-1}_{i,j}$ at time t1t-1 and $u^{t}_{i,j}$ at time $t$, and the one-step prediction $\hat{u}^{t+1}_{i,j}$ at time $t+1$, the discretized PDE residual is represented as:
@@
Hence, $\frac{\partial L_{\text{PDE}}}{\partial\hat{u}^{t+1}}$ does not depend on the prediciton $\hat{u}^{t+1}$ itself. Notably, for the wave equation, being a linear system, the PDE residual in Equation 3.1 does not rely on the assumption of small prediction error, as no Taylor expansion is needed. This further implies that PhysicsCorrect remains robust to prediction errors in linear systems, a phenomenon that will be demonstrated empirically in later sections.
Kuramoto-Sivashinsky Equation.
The Kuramoto-Sivashinsky equation with a constant viscosity parameter of 1.0 is as follows:
@@
Heuristic Theoretical Analysis.
We provide a heuristic theoretical analysis to offer intuitive insight into the behavior of the proposed method (A schematic diagram is shown in Figure 9). Under the linearized correction, let $r=L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1})$ denote the current PDE residual, and $r^{+}=L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1}+\mathbf{u}_{c})$ the updated residual after applying the correction $\mathbf{u}_{c}$. Linearizing around $\hat{\mathbf{u}}_{t+1}$ gives
$r^{+}\approx r+A\mathbf{u}_{c},\qquad A=\frac{\partial L_{\text{PDE}}}{\partial\hat{\mathbf{u}}_{t+1}}.$ (15)
Solving it in the least-squares sense yields
$\mathbf{u}_{c}^{*}=-A^{\dagger}r,\quad r^{+}\approx(I-AA^{\dagger})r,$ (16)
where $A^{\dagger}$ is the Moore–Penrose pseudoinverse and $AA^{\dagger}$ the orthogonal projector onto $\mathrm{range}(A)$. Each correction step therefore removes the component of the residual lying in $\mathrm{range}(A)$, strictly reducing its norm unless $r\perp\mathrm{range}(A)$. This shows that the correction projects predictions back onto the PDE-consistent manifold, mitigating error accumulation in long rollouts. With a relaxed update $\mathbf{u}_{c}=-\gamma A^{\dagger}r$ ($0<\gamma<2$), under which
$r^{+}=(I-\gamma AA^{\dagger})r$
acts as a contraction on $\mathrm{range}(A)$, ensuring residual decay and connecting the theoretical bound to our empirical stability observations.
Figure 9: Schematic illustration of the PhysicsCorrect predictor–corrector workflow. Starting from the true trajectory $(\mathbf{u}_{0},\dots,\mathbf{u}_{t})$, a pretrained neural PDE solver produces baseline predictions $(\hat{\mathbf{u}}_{1},\dots,\hat{\mathbf{u}}_{t})$, which gradually deviate from the numerical reference. Our method applies a physics-informed correction at each rollout step, projecting $\hat{\mathbf{u}}_{t}$ back toward the PDE-consistent manifold and preventing long-term error accumulation.
@@ Appendix B Experimental Details
In this section, we introduce the details of the configurations of different benchmarks and neural network configurations, as well as additional visualization results.
@@ B.1 Navier-Stokes Equation
Regarding the neural network architectures, we implement the standard Fourier Neural Operator (FNO) following (Li et al. 2020b), which consists of four Fourier layers. Each layer uses 12 Fourier modes in both the $x$ and $z$ directions, and the width of the feature map is set to 64. The activation function used is the Gaussian Error Linear Unit (GELU) (Hendrycks and Gimpel 2016). After the Fourier layers, a linear projection is applied to map the high-dimensional latent representation back to the spatial domain. For the UNet architecture, we adopt the configuration described in (Huang et al. 2023), with a hidden size of 128. For the Vision Transformer (ViT), we use a patch size of 8 for spatial embedding and stack 4 standard pre-norm transformer blocks. Each block contains four attention heads, with each head having a dimensionality of 128. GELU is also used as the activation function for ViT.
@@
We fixed the wave speed $c$ as 1.0 in our test. We generate Gaussian random fields as initial conditions and then use fourth-order central differences for spatial derivatives and a fourth-order Runge–Kutta (RK4) scheme for time integration, to accurately evolve the displacement field, following (Rosofsky, Majed, and Huerta 2022). The data are generated on a $\times128$ grid with a time step of 1e-4 and recorded every 100 time steps. We sampled 512 trajectories with the first 10 steps (with an interval of 1e-2) for training and another 64 trajectories with all 100 steps for testing. We use an initial learning rate of 1e-3 with a decay factor of 0.6 every 100 epochs.
@@
First-order residual vs second-order residual.
@@ B.3 Kuranmoto-Sivashinksy Equation
We aim to solve the Equation 13 on a temporal domain $[0,100]$ and a spatial domain $[0,64]$ while focusing on the chaotic part. The equation is numerically solved using the spectral method (Dresdner et al. 2023), with a spatial interval of 0.125 and a temporal interval of 0.05, for data generation. We begin from $v(x,50)$ and train a one-step predictor using neural networks to learn the residual $\delta v^{t}=v^{t+1}-v^{t}$. The training set includes 512 trajectories, using only the first 500 time steps of each, while the test set contains 64 additional trajectories, each with the full 1,000-step rollout. All models are trained with a learning rate of 1e-3, a batch size of 256, and a total of 3,000 training epochs.
For the network configurations, we use a one-dimensional variant of the FNO (FNO-1D), adapted from its 2D counterpart by replacing all 2D convolutional operations with 1D convolutions. The number of Fourier modes is set to 16, and the feature map width is 64. GELU is used as the activation function. The 1D UNet is similarly adapted from the 2D UNet by substituting all 2D convolutional layers with 1D versions. The hidden size is set to 64. We use the LeakyReLU activation function with a negative slope of 0.1, along with Group Normalization (Wu and He 2018). For the 1D ViT, we use a patch size of 4 and an embedding dimension of 768. The model consists of 4 transformer blocks, each containing 16 attention heads with a head dimension of 768. GELU is employed as the activation function throughout the model.
@@
While chaotic PDEs such as the KS equation exhibit extreme sensitivity to initial conditions, their trajectories remain confined to a low-dimensional manifold defined by the governing equations. Our correction step, even when based on an approximate and fixed Jacobian, serves as a coarse projection onto this manifold. This projection regularizes the neural predictions by suppressing unstable high-frequency modes and redirects the state evolution toward dynamically consistent directions, preventing accumulation of off-manifold drift. Therefore, perfect inversion of the true Jacobian is less critical than accurate residual definition. As long as the correction step correlates with the true manifold-normal direction, stability can be maintained. This explains why our cached, approximate Jacobian yields substantial improvements even in chaotic regimes.
B.4 Comparison with other baselines
We compared PhysicsCorrect with PDE-Refiner (Lippe et al. 2023). Under 1-D KS settings, PhysicsCorrect achieves over 3× lower cost, 3× faster inference, while reducing the 1000-step rollout error from 1.892 (PDE-refiner) to 0.461 (our approach), highlighting the efficiency of our method. PhysicsCorrect is orthogonal to training-based methods (e.g., DPOT (Hao et al. 2024)) and can be seamlessly integrated with them.
B.5 Additional Implementation details
Compute Resources.
All training is performed on a single NVIDIA A6000 GPU. For the Navier-Stokes equation, training for 30,000 steps with a batch size of 64 takes roughly 60, 12, and 170 minutes using FNO, U-Net, and ViT, respectively. For the wave equation, training for 1,000 epochs takes roughly 29, 6, and 65 minutes using FNO, U-Net, and ViT, respectively. For the KS equation, training for 3,000 epochs takes roughly 100 minutes using FNO, 70 minutes using U-Net, and approximately 5 hours using ViT.
Negligible Additional Inference Time.
We also evaluate the inference time of both the baseline models and their counterparts augmented with our proposed correctors. As shown in Figure 4, the ratio between the inference time of the baseline with corrector and that of the baseline model is close to 1.0. Although the Navier–Stokes case with the ViT baseline shows a slightly higher ratio, due to the lightweight nature of the baseline, the additional inference time introduced by our corrector remains negligible, particularly for larger-scale problems and more complex neural networks.
B.6 Sensitivity Analysis: Error Structure and Magnitude Limits